# Transfer of quantum game strategies

### v1

@@ 6.4. An algebraic approach
A quantum non-local game $\varphi:\mathcal{P}_{XX}\rightarrow\mathcal{P}_{AA}$ is concurrent if $\varphi(\tilde{J}_{X})=\tilde{J}_{A}$. A QNS correlation $\Phi\in\mathcal{Q}_{\rm ns}$ is called concurrent if $\Phi(\tilde{J}_{X})=\tilde{J}_{A}$; equivalently, if $\Phi$ is a perfect strategy for the (trivial) implication game $\varphi_{\tilde{J}_{X}\rightarrow\tilde{J}_{A}}$. The study of concurrent games and their perfect strategies has grown in recent years (see [6, 7, 38]), and is of particular interest for its connections to the study of quantum automorphism groups and compact quantum groups, along with its ramifications for the study of synchronous games; indeed— concurrent games were introduced as a “quantization” of synchronous games. The reasoning behind this claim comes from [6, Remark 2.1]: for classical non-local game $\lambda:X\times X\times A\times A\rightarrow\{0,1\}$, then $\lambda$ is synchronous if and only if $\varphi_{\lambda}(J_{X}^{\rm cl})\leq J_{A}^{\rm cl}$. Concurrency allows us to see how a correlation affects the “non-classical” portion of the (unnormalized) maximally entangled state- ideally, sending it to a state supported on $\tilde{J}_{A}$.
@@
(33) $\sum\limits_{x_{2},x_{2}^{\prime}\in X_{2}}\sum\limits_{x_{1},x_{1}^{\prime}\in X_{1}}e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\leq|X_{2}||X_{1}|1.$
@@
(34) $\tilde{s}(h_{x_{2}x_{2}^{\prime},x_{1}x_{1}}^{*}h_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}})=0,\;\;\;\;x_{i},x_{i}^{\prime}\in X_{i},i=1,2.$
@@
(35) $ze_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\otimes 1\otimes v\sim z\otimes e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}\otimes v\sim e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}z\otimes 1\otimes v,$
@@
(36) $u\otimes ze_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes 1\sim u\otimes z\otimes e_{a_{1}^{\prime}a_{1},a_{2}^{\prime}a_{2}}\sim u\otimes e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}},$
@@
(37) $ze_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\otimes 1\otimes z^{\prime}e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes 1\sim z\otimes e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}\otimes z^{\prime}\otimes e_{a_{1}^{\prime}a_{1},a_{2}^{\prime}a_{2}}$
(38) $\sim e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}z\otimes 1\otimes e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}z^{\prime}\otimes 1,$
@@
(ii) If $\Gamma$ is an approximately quantum jointly tracial correlation, again by the proof of Theorem 4.12 there exists a state $s:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\rightarrow\mathbb{C}$ such that $\Gamma=\Gamma_{s}$. As each approximately quantum SQNS correlation is quantum commuting, by (i) there exists some trace $\mathsf{T}:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\rightarrow\mathbb{C}$ such that $\Gamma=\Gamma_{\mathsf{T}}$ as well. By [38, Lemma 9.2], for any finite sets $X,Y$ there exists a *-isomorphism $\partial:\mathcal{C}_{X,Y}\rightarrow\mathcal{C}_{X,Y}^{\rm op}$ such that
@@
Let $\partial_{X}:\mathcal{C}_{X_{2},X_{1}}\rightarrow\mathcal{C}_{X_{2},X_{1}}^{\rm op}$ and $\partial_{A}:\mathcal{C}_{A_{1},A_{2}}\rightarrow\mathcal{C}_{A_{1},A_{2}}^{\rm op}$ be the *-isomorphisms corresponding to each ${\rm C}^{*}$-algebra. Let
@@
(39) $\tilde{s}(u)=\langle(\pi_{X}\otimes\pi_{X^{\prime}}\otimes\pi_{A}\otimes\pi_{A^{\prime}})(u))\xi,\xi\rangle,$
@@
That this is a state follows from [37, Proposition 4.23]; that it is also tracial follows from a standard argument using the (automatic) continuity of the mapping on the maximal tensor product, in conjunction with calculations on the dense subset of finite linear combinations of simple tensors. Define tracial state $\tilde{\tau}$ on $\mathcal{C}_{X_{2},A_{2}}$ in the following way: by Remark 6.12, there exists a surjective *-representation $\pi:\mathcal{C}_{X_{2},A_{2}}\rightarrow\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{X_{1},A_{1}}$ sending $\pi(e_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}})=\tilde{e}_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}}$ for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$. Using $\pi$, we let
(40) $\tilde{\tau}(u):=((\mathsf{T}\odot\tau)\circ\pi)(u),\;\;\;\;u\in\mathcal{C}_{X_{2},A_{2}}.$
@@
(41) $\gamma_{P,Q}:M_{XX}\otimes M_{AA}\otimes\mathfrak{C}_{X,A}\otimes\mathfrak{C}_{X,A}^{\rm op}\rightarrow\mathfrak{C}_{X,A}$
@@
be the generated *-ideal in $\mathfrak{C}_{X,A}$, and $J(\varphi)$ be the corresponding ideal considered in $\mathcal{C}_{X,A}$. Finally, write $\mathfrak{C}(\varphi)=\mathfrak{C}_{X,A}/\mathfrak{J}(\varphi)$ (resp. $\mathcal{C}(\varphi)=\mathcal{C}_{X,A}/J(\varphi)$) for the quotient *-algebra (resp. quotient ${\rm C}^{*}$-algebra). Perfect strategies for a concurrent quantum game $\varphi$ were shown to correspond to tracial states acting on *-representations of $\mathcal{C}_{X,A}$ which annihilate $\mathfrak{J}(\varphi)$ or $J(\varphi)$ in [6, 7]. In light of previous results, we may conclude by giving an algebraic characterization of our transfer of perfect strategies between quantum games.
Theorem 6.18.
@@
Recall, that, for finite set $X$, the ${\rm C}^{*}$-algebra of the free unitary quantum group $C(\mathcal{U}_{X}^{+})$ is the universal unital ${\rm C}^{*}$-algebra generated by the entries $u_{x,a}$ of an $X\times X$ matrix $U=(u_{x,a})_{x,a\in X}$ under the condition that $U$ and $U^{\rm t}$ are unitary (known otherwise as a bi-unitary matrix). Subalgebra $C(\mathbb{P}\mathcal{U}_{X}^{+})$ is generated by length two words of the form $u_{x,a}^{*}u_{x^{\prime},a^{\prime}},x,x^{\prime},a,a^{\prime}\in X$- and was shown in [1, Corollary 4.1] to be the proper quantization of the automorphism group of $M_{X}$. In [7, Theorem 6.7], concurrent bicorrelations of types ${\rm t}\in\{\rm loc,q,qc\}$ (and therefore, strategies for the quantum graph isomorphism game) were shown to be in correspondence with different tracial states on $C(\mathbb{P}\mathcal{U}_{X}^{+})$.
@@
Finally, set $\mathcal{A}_{P,P}:=C(\mathbb{P}\mathcal{U}_{X}^{+})/\mathcal{I}_{P,P}$. In [7, Remark 7.12], it was shown that ${\rm C}^{*}$-algebra $\mathcal{A}_{P,P}$ can be endowed with a natural co-associative comultiplication $\Delta_{P}:\mathcal{A}_{P,P}\rightarrow\mathcal{A}_{P,P}\otimes\mathcal{A}_{P,P}$; this means $\mathcal{A}_{P,P}$ can be viewed as a compact quantum group. By [7, Theorem 7.10] and Theorem 6.9, we see that the existence of a quantum hypergraph isomorphism $\Gamma$ between $\overline{\mathcal{U}}_{P_{1},P_{1}}$ and $\mathcal{U}_{P_{2},P_{2}}$ ensures a way to construct *-representations of compact quantum group $\mathcal{A}_{P_{2},P_{2}}$ from *-reps of $\mathcal{A}_{P_{1},P_{1}}$; that is, a way to transfer quantum automorphisms between quantum graphs.

### v2

@@ 6.2. Concurrent strategies
The study of concurrent games and their perfect strategies has grown in recent years (see [6, 7, 39]), and is of particular interest for its connections to the study of quantum automorphism groups and compact quantum groups, along with its ramifications for the study of synchronous games; indeed— concurrent games were introduced as a “quantization” of synchronous games. The reasoning behind this claim comes from [6, Remark 2.1]: for classical non-local game $\lambda:X\times X\times A\times A\rightarrow\{0,1\}$, then $\lambda$ is synchronous if and only if $\varphi_{\lambda}(J_{X}^{\rm cl})\leq J_{A}^{\rm cl}$. Concurrency allows us to see how a correlation affects the “non-classical” portion of the (unnormalized) maximally entangled state- ideally, sending it to a state supported on $\tilde{J}_{A}$.
@@
Let $\Gamma:M_{X_{2}X_{2}\times A_{1}A_{1}}\rightarrow M_{X_{1}X_{1}\times A_{2}A_{2}}$ be an SQNS correlation, and let $\sigma\in M_{A_{1}A_{1}}$ be an arbitrary state. Define the mapping $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}:M_{X_{2}X_{2}}\rightarrow M_{X_{1}X_{1}}$ via
One easily shows that $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}$ is a well-defined QNS correlation; furthermore, $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma}=\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}^{\sigma^{\prime}}$ for any states $\sigma,\sigma^{\prime}\in M_{A_{1}A_{1}}$. Thus, we let $\Gamma_{X_{2}X_{2}\rightarrow X_{1}X_{1}}$ denote the right marginal channel of $\Gamma$, for an arbitrary fixed state $\sigma\in M_{A_{1}A_{1}}$. An analogous argument shows that $\Gamma^{A_{1}A_{1}\rightarrow A_{2}A_{2}}$ is a well-defined left marginal QNS correlation for $\Gamma$, for an arbitrary fixed state $\rho\in M_{X_{2}X_{2}}$.
Remark 6.14.
Note that, in the event $\Gamma\in\mathcal{C}_{\rm st}$ for ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$, then the marginal channels previously discussed (when considering $\Gamma$ inside $\mathcal{Q}_{\rm st}$) coincide with the marginal channels discussed in [20, Section 5].
@@
Remark 6.16.
There exists a surjective *-homomorphism $\pi:\mathcal{C}_{X_{2},A_{2}}\rightarrow\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{X_{1},A_{1}}$. Indeed: for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$ let
@@
$\sum\limits_{x_{2},x_{2}^{\prime}\in X_{2}}\sum\limits_{x_{1},x_{1}^{\prime}\in X_{1}}e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\leq|X_{2}||X_{1}|1.$ (28)
@@
$\tilde{s}(h_{x_{2}x_{2}^{\prime},x_{1}x_{1}}^{*}h_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}})=0,\;\;\;\;x_{i},x_{i}^{\prime}\in X_{i},i=1,2.$ (29)
@@
$ze_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\otimes 1\otimes v\sim z\otimes e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}\otimes v\sim e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}z\otimes 1\otimes v,$ (30)
@@
$u\otimes ze_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes 1\sim u\otimes z\otimes e_{a_{1}^{\prime}a_{1},a_{2}^{\prime}a_{2}}\sim u\otimes e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}},$ (31)
@@
$ze_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}\otimes 1\otimes z^{\prime}e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes 1\sim z\otimes e_{x_{2}^{\prime}x_{2},x_{1}^{\prime}x_{1}}\otimes z^{\prime}\otimes e_{a_{1}^{\prime}a_{1},a_{2}^{\prime}a_{2}}$ (32)
$\sim e_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}z\otimes 1\otimes e_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}z^{\prime}\otimes 1,$ (33)
@@
(ii) If $\Gamma$ is an approximately quantum jointly tracial correlation, again by the proof of Theorem 6.6 there exists a state $s:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\rightarrow\mathbb{C}$ such that $\Gamma=\Gamma_{s}$. As each approximately quantum SQNS correlation is quantum commuting, by (i) there exists some trace $\mathsf{T}:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\rightarrow\mathbb{C}$ such that $\Gamma=\Gamma_{\mathsf{T}}$ as well. By [39, Lemma 9.2], for any finite sets $X,Y$ there exists a *-isomorphism $\partial:\mathcal{C}_{X,Y}\rightarrow\mathcal{C}_{X,Y}^{\rm op}$ such that
@@
[…]
@@
$\tilde{s}(u)=\langle(\pi_{X}\otimes\pi_{X^{\prime}}\otimes\pi_{A}\otimes\pi_{A^{\prime}})(u))\xi,\xi\rangle,$ (34)
@@
That this is a state follows from [38, Proposition 4.23]; that it is also tracial follows from a standard argument using the (automatic) continuity of the mapping on the maximal tensor product, in conjunction with calculations on the dense subset of finite linear combinations of simple tensors. Define tracial state $\tilde{\tau}$ on $\mathcal{C}_{X_{2},A_{2}}$ in the following way: by Remark 6.16, there exists a surjective *-representation $\pi:\mathcal{C}_{X_{2},A_{2}}\rightarrow\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{X_{1},A_{1}}$ sending $\pi(e_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}})=\tilde{e}_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}}$ for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$. Using $\pi$, we let
$\tilde{\tau}(u):=((\mathsf{T}\odot\tau)\circ\pi)(u),\;\;\;\;u\in\mathcal{C}_{X_{2},A_{2}}.$ (35)
@@
$\gamma_{P,Q}:M_{XX}\otimes M_{AA}\otimes\mathfrak{C}_{X,A}\otimes\mathfrak{C}_{X,A}^{\rm op}\rightarrow\mathfrak{C}_{X,A}$ (36)
@@
[…]
Theorem 6.19
@@
[…]
@@
[…]